1. Ralph owns a small pizza restaurant, where he works full-time in the kitchen. His total revenue last year was $100,000, and his rent was $3,000 per month. He pays his one employee $2,000 per month, and the cost of ingredients and overhead averages $500 per month. Ralph could earn $35,000 per year as the manager of a competing pizza restaurant nearby. His total explicit costs for the year were

a. $24,000.

b. $66,000.

c. $72,000.

d. $60,000.

e. $6,000.

2. Ralph owns a small pizza restaurant, where he works full-time in the kitchen. His total revenue last year was $100,000, and his rent was $3,000 per month. He pays his one employee $2,000 per month, and the cost of ingredients and overhead averages $500 per month. Ralph could earn $35,000 per year as the manager of a competing pizza restaurant nearby. His total accounting profit for the year was:

a. –$1,000.

b. $34,000.

c. $35,000.

d. $100,000.

e. $72,000.

3. Ralph owns a small pizza restaurant, where he works full-time in the kitchen. His total revenue last year was $100,000, and his rent was $3,000 per month. He pays his one employee $2,000 per month, and the cost of ingredients and overhead averages $500 per month. Ralph could earn $35,000 per year as the manager of a competing pizza restaurant nearby. His total economic profit for the year was:

a. –$35,000.

b. –$1,000.

c. $20,000.

d. $34,000.

e. $65,000.

4. Lauren owns a bakery that produces, among other things, wedding cakes. She currently has 5 employees; with 5 employees, her bakery can produce 7 wedding cakes per day. If she hired a sixth employee, she’d be able to produce 9 wedding cakes per day. Therefore, the marginal product of the sixth employee is __________ wedding cake(s).

a. 2

b. 5

c. 9

d. 7

e. 1.5

5. Use the following scenario to answer the questions that follow.
Steve owns a bike store. His total costs are $1.2 million per year, his variable costs are $750,000, and his fixed costs are $450,000 per year. Last year, Steve sold 1,200 bikes.

a. Steve’s average total cost was __________ per bike.

b. Steve’s average fixed cost was __________ per bike.
Answer #1

1. b. $66,000
(Total explicit cost for the year = total direct costs for the year = (3000 + 2000 + 500)*12 = 5500*12 = 66000)

2. b. $34,000.
(Total accounting profit = total revenue - total explicit cost = 100,000 - 66000 = 34,000)

3. b. –$1,000
(Total economic profit = total revenue - total explicit cost - implicit cost = 100,000 - 66000 - 35,000 = -1,000)

4. a. 2
(Marginal product of 6th employee = Change in total output = 9-7 = 2)

5. a. Average total cost = Total cost/Quantity = 1.2 million/1200 = 1,200,000/1200 = 1000 per bike
b. Average fixed cost = Total fixed cost/Quantity = 450,000/1200 = 375 per bike